# Strong Coupling Constant Determination from the new CTEQ-TEA Global QCD Analysis

### v1

@@ II.2 The CTEQ-TEA goodness-of-fit function
where $\chi^{2}_{R}$ quantifies the agreement of theory with experiments, and $\chi^{2}_{\textrm{LM}}$ is a generally small contribution imposing theoretical priors on the PDFs. The latter is constructed from Lagrange-Multiplier (LM) contributions whose role is to (i) enforce physical behavior of PDF flavor ratios in the extrapolation regions with poor constraints from the data and (ii) enable a “ridge” (L2) regularization on a few poorly constrained, yet correlated PDF parameters.
@@
$\chi^{2}_{E}(\bm{a},\bm{\lambda},\alpha_{s}(M_{Z}),\{m_{q}\})=$
$\sum_{k=1}^{N_{\textrm{pt}}^{(E)}}\frac{1}{s_{k}^{2}}\left(D_{k}-T_{k}\left(\bm{a},\alpha_{s}(M_{Z}),\{m_{q}\}\right)-\sum_{\alpha=1}^{N_{\lambda}^{(E)}}\beta_{k\alpha}\left(\bm{a},\alpha_{s}(M_{Z}),{m_{q}}\right)\lambda_{\alpha}\right)^{2}+\sum_{\alpha=1}^{N_{\lambda}^{(E)}}\lambda_{\alpha}^{2}\,.$ (3)
Here $D_{k}$ and $T_{k}$ are the $k$-th central experimental and theory values, respectively, $s_{k}=\sqrt{s^{2}_{k,\textrm{stat}}+s^{2}_{k,\textrm{uncor sys}}}$ is the total uncorrelated error on the measurement $D_{k}$, and $\beta_{k\alpha}^{(E)}$ is the correlation matrix of systematic uncertainties. $\bm{a}\equiv\{a_{1},a_{2}\dots,a_{N_{\textrm{pdf}}}\}$ and $\bm{\lambda}\equiv\{\lambda_{1},...\lambda_{N_{\lambda}^{(E)}}\}$ are the vectors of PDF parameters and systematic nuisance parameters of $E$, with the latter commonly assumed to be random and distributed according to a standard normal distribution $\mathcal{N}(0,1)$. Equation (3) indicates that $\chi^{2}$ generally depends on QCD parameters – $\alpha_{s}(M_{Z})$ and quark masses $m_{q}$ – both through the theoretical predictions $T$ and the experimental correlation matrices $\beta_{k\alpha}$, although this dependence of the latter is usually neglected.
In our study, we minimize the total $\chi^{2}$ with respect to $\alpha_{s}(M_{Z})$ and $\bm{a}$, while using constant $\{m_{q}\}$ and profiling [55] the nuisance parameters $\bm{\lambda}$ for every $\{\alpha_{s},\bm{a}\}$ combination. We will find it helpful that, at its global minimum for each $\alpha_{s}(M_{Z})$, the $\chi^{2}$ can be expressed in terms of the best-fit PDF parameter combination $\bm{a}_{0}$ as (see Eq. (B8) in [16])
$\min_{\bm{a}}\chi^{2}(\bm{a},\alpha_{s}(M_{Z}),\{m_{q}\})\equiv\chi^{2}(\bm{a}_{0},\alpha_{s}(M_{Z}),\{m_{q}\})=\sum_{E}\left[\underbrace{\sum_{k=1}^{N_{\textrm{pt}}^{(E)}}[r_{0,k}^{(E)}]^{2}}_{\equiv D^{2}}+\underbrace{\sum_{\alpha=1}^{N_{\lambda}^{(E)}}[\lambda^{(E)}_{0,\alpha}]^{2}}_{\equiv R^{2}}\right],$ (4)
@@
and all terms on the right-hand side corresponding to experiment $E$. Furthermore, since $r_{0,k}^{(E)}$ and $\lambda^{(E)}_{0,\alpha}$ are correlated only through a relatively small number of PDF parameters, especially across separate experiments, in the first approximation they all can be assumed to be independently and normally distributed, $r_{0,k}^{(E)}\sim\mathcal{N}(0,1)$ and $\lambda^{(E)}_{0,\alpha}\sim\mathcal{N}(0,1)$. This independence opens the possibility of reclustering the individual data points in the total $\chi^{2}$ in Eq. (4), which will be pursued in Sec. IV. Furthermore, in Sec. IV.4 we can examine separate reclustering of sums of all best-fit residuals and nuisance parameters, denoted as $D^{2}$ and $R^{2}$, respectively.
II.3 Limitations of exisiting models for correlated systematics
Several aspects of the implementation of correlated systematic errors using $\beta_{k\alpha}$ in our $\chi^{2}$ will be relevant for the interpretation of the results that will follow.
@@ II.3 Limitations of exisiting models for correlated systematics
    We first note that the total number $N_{\lambda}$ of the experimental nuisance parameters has grown from at most a few tens in the early global fits to more than 2000 in the CT25 analysis. That is, $N_{\lambda}$ is now close to a half of the total data points $N_{\textrm{pt}}$ due to the tendency of the latest experiments to provide large correlation or covariance matrices.
@@
    While the common uncertainty estimates in the PDF fits operate in the traditional central limit applicable when the number of data, $N_{\textrm{pt}}$, is much larger than the number of parameters $N_{\textrm{par}}$, the reality of the latest fit is that its total $N_{\textrm{par}}=N_{\textrm{pdf}}+N_{\lambda}$ and $N_{\textrm{pt}}$ are comparable. The central limit estimates may receive corrections of order $N_{\textrm{par}}/N_{\textrm{pt}}$, which are currently not considered in the PDF analyses.
@@
    Considerable attention has been recently paid to the fact that, by using the linear Gaussian models to publish their correlated systematic uncertainties, experiments unavoidably leave significant information out of the scope accessible within the PDF fit. There are a couple such points, first, related to the standard assertion that all systematic nuisance parameters follow the standard normal distribution, $\lambda^{(E)}_{\alpha}\sim\mathcal{N}(0,1)$ in the $\chi^{2}$ definition in Eq. (3). Not only it is highly unlikely that all $\lambda_{\alpha}$ parameters will be thusly distributed, there are known important cases when this assumption is false, e.g., for so-called two-point uncertainties that take the difference of at most two estimates, for instance, provided by two parton shower programs, as a measure of a possible uncertainty.
  4. 4.
    The second point of the information loss emerges from the common practice of publishing only normalized correlation matrices, $\sigma_{k\alpha}$, from which the actual correlation matrices, $\beta_{k\alpha}=X_{k}\sigma_{k\alpha}$ in Eq. (3), must be reconstructed by multiplying by unspecified reference values $X_{k}$ that are chosen within the PDF fit itself. The $X_{k}$ values are of order of either the central data values, $D_{k}$, or respective theoretical values, $T_{k}$. They don’t need to coincide with either; the prescription for choosing $X_{k}$ is a delicate part of the fitting model [56, 57], approached differently by each PDF fitting group, who try to reasonably fill the lacunas in the correlation models they get.
How do these considerations impact the precise determination of the QCD coupling constant? We address this question in the next section, after discussing another potential source of uncertainty due to the freedom in the PDF parametrization.
@@ III.1 The PDF parametrization uncertainty is small…
It has been pointed out over years, e.g., in  [15, 16, 58, 59], that QCD parameters determined from a global fit are mutually dependent with the PDFs and thus must be fitted simultaneously with the PDFs. Within the global fit, the correlation between the $\alpha_{s}(M_{Z})$ and PDFs can be captured either by fitting them together or by scanning the $\chi^{2}(\alpha_{s}(M_{Z}),\vec{a}_{0})$ over $\alpha_{s}(M_{Z})$, while minimizing the $\chi^{2}$ with respect to the PDFs for each $\alpha_{s}(M_{Z})$ value. Either approach can be used to determine the best-fit combination of $\alpha_{s}(M_{Z})$ and PDFs together with their uncertainties. We will employ both approaches, especially the latter one – the $\alpha_{s}$ scan – to elucidate such combined fits.
Our study suggests a low correlation of the best-fit $\alpha_{s}(M_{Z})$ with the functional forms of PDF parametrizations in the vicinity of the overall global $\chi^{2}$ minimum. The consequence is that the choice of the PDF parametrizations has a low impact on the $\alpha_{s}(M_{Z})$ and its error, among the candidate functional forms of the CT25 fit that render the lowest achieved values of $\chi^{2}$. The variation of the best-fit $\alpha_{s}(M_{Z})$ caused by the choice of the parametrization forms is comparable with the uncertainty corresponding to a $\Delta\chi^{2}=1$ change in the global $\chi^{2}$. We will see that this is quite a small uncertainty, compared to other sources.
@@
Figure 1 illustrates the low correlation of best-fit $\alpha_{s}(M_{Z})$ and PDFs by comparing the $\chi^{2}$ dependence from an $\alpha_{s}(M_{Z})$ scan for the functional forms of the nominal CT25 fit and 287 fits with alternative PDF parametrizations explored as a part of the full CT25 exercise. In this plot, we focus on the component of the total $\chi^{2}$ consisting from experimental log-likelihoods and denoted by $\chi^{2}_{R}$ as in Eq. (1), and neglecting the prior component $\chi^{2}_{\textrm{LM}}$ that is relatively constant in this comparison. All fits in this figure are done with identical settings that are close both to the CT25prel and final CT25 ones introduced in Sec. II.1. Namely, the shown fit with the nominal CT25 parametrization achieves $\min\chi^{2}_{R}\equiv\chi^{2}_{0}=5087$ for $N_{\textrm{pt}}=4260$ data points. Compared to the CT25prel fit, it excludes the CDHSW heavy-nucleus CC DIS data set [48], which prefers a low $\alpha_{s}(M_{Z})$, as well as the weakly sensitive H1 $F_{L}$ data set [52]. It retains other experiments of the CT25prel baseline.
The orange star in Fig. 1 indicates the best-fit combination $\{\alpha_{s}\approx 0.1185,\chi^{2}_{R}=5087\}$ obtained with the nominal CT25 parametrization. The approximately parabolic orange curve indicates the $\chi^{2}_{R}$ dependence from the $\alpha_{s}$ scan with this parametrization. The vertical magenta band indicates the interval of $\alpha_{s}(M_{Z})$ values corresponding to the increase $\Delta\chi^{2}_{R}\leq 1$ from the $\alpha_{s}$ scan, measured with respect to the minimal $\chi^{2}_{R}$ on this quasi-parabola.
@@
Here we observe a striking behavior: while the $\chi^{2}$ variations among the candidate fits are sizable in the vertical ($\chi^{2}$) direction, they have modest horizontal variations, compatible with the size of the $\Delta\chi^{2}=1$ interval, as far as the best-fit $\alpha_{s}(M_{Z})$ are concerned. The picture immediately reminds of the textbook case of the $\chi^{2}$ variability due to approximately Gaussian random fluctuations in data [60]: when one plots the log-likelihood ($\chi^{2}$) vs. the parameter of interest ($a$), vertical fluctuations of the resulting parabola fall into the interval $\chi^{2}_{R}\leq\chi^{2}_{0}+\sqrt{2N_{\textrm{pt}}}$ in $\approx 68\%$ of cases, while about 68% of the horizontal fluctuations fall within the interval $\Delta\chi^{2}_{R}\leq 1$.
In our example, the best-fit $\alpha_{s}(M_{Z})$ from fits with alternative parametrizations fall within the interval for $\Delta\chi^{2}_{R}=1$ from the scan, reflecting statistical fluctuations in the data at 68% CL for the fixed data set and PDF functional forms. This interval corresponds to the uncertainty $\delta\alpha_{s}\approx\pm 0.0005$, and this conclusion holds for other candidate selections of baseline experiments. Namely, while the preferred $\alpha_{s}(M_{Z})$ value moves between $\approx 0.1175$ and 0.1185 for the CT25prel and CT25 selections of data sets, see the bottom of Table 1, the pattern of the parametrization dependence is about the same in these fits as in Fig. 1, with the parametrization uncertainty staying at the level of 0.0005.
The conclusion based on the ensemble of nearly 300 PDF parametrizations is then that they give very similar $\alpha_{s}(M_{Z})$ values, showing that the determination of $\alpha_{s}(M_{Z})$ is essentially independent of the chosen PDF parametrization. If only this source of uncertainty were considered, the uncertainty span of approximately $\pm 0.0005$ about the central value would be smaller than the PDG uncertainty [61] of $\pm 0.0009$ and even dedicated PDF-only fits [62, 17, 63, 20, 19, 16, 64] quoting errors of order $\pm 0.002$.
@@
The global fit contains two more significant sources of $\alpha_{s}$ variations, introduced by the dependence on the models of correlated systematic errors already brought up in Sec. II.3, as well as by the more-than-normal disagreements between the parts of the hadronic data set. This section will address the first aspect and illustrate the possible impact of the choice of the reference normalization factors $X_{k}$ in the experimental correlation matrices $\beta_{k\alpha}=X_{k}\sigma_{k\alpha}$, which is controlled in our fit by the parameter ksys. Section IV will follow with addressing the second aspect.
Since the experiments normally do not publish their recommended $X_{k}$, a variety of “guesses” about $X_{k}$ can be, and have been, made in the past PDF fits [56, 57]. As illustrative possibilities, we consider two such prescriptions, which we label as “ksys=1” and “ksys=2”. According to the ksys=1 prescription, each correlated systematic element $\beta_{k\alpha}$ (initially provided as a percentage fraction $\sigma_{k\alpha}$) is normalized to the current theory prediction $T_{k}$ for each data point. This treatment is more suitable when the corresponding experimental error is multiplicative. On the other hand, it can potentially introduce a theory bias in the effect of systematic errors. The majority of fits by CTEQ-TEA and other groups choose ksys=1 for all correlated errors as a default.
@@
For the ensuing discussion, we divide our baseline data set of measurements into three subsets: 1) deep-inelastic scattering (DIS), 2) Drell-Yan (DY) pair production, 3) inclusive jet and $t\bar{t}$ production. In each subset, we normalize all $\beta_{k\alpha}$ as though they are multiplicative (ksys=1) or additive (ksys=2). In reality, most measurements contain a mix of multiplicative and additive errors. The true value of $\alpha_{s}(M_{Z})$ extracted from the global fit is somewhere between the extremes corresponding to the “all ksys=1” and “all ksys=2” options. This setup allows us to explore sensitivity to systematic error modeling, particularly for hadron jet production observables that exhibit a very large sensitivity.
We first demonstrate what happens in the final CT25 fit when we change the default setting of using “ksys=1” in all three measurement categories, labeled as $\{\texttt{ksys}_{DIS},\texttt{ksys}_{DY},\texttt{ksys}_{\textrm{Jet}+t{\bar{t}}}\}$ = $\{1,1,1\}$. First, to illustrate how the ksys choice modifies $\chi^{2}$ when all other settings are unchanged, let us examine the $\chi^{2}$ dependence for several combinations of ksys values while using the same PDF ensemble. Figure 2 illustrates these changes, by showing $\chi^{2}$ vs $\alpha_{s}$ using the nominal CT25 data set and best-fit PDFs for $\{1,1,1\}$, while computing the shown $\chi^{2}$ respectively with $\{1,1,1\}$, $\{1,1,2\}$, $\{1,2,2\}$, and $\{2,2,2\}$ ksys choices.
@@
It is interesting to delve into details of the fits to understand the driving mechnanism of these shifts. First, it appears that the shifts are largely driven by the decrease in the average of data point residuals and less by the best-fit nuisance parameters. This can be seen from Fig. 3 that separately plots the contributions from the best-fit residuals and nuisance parameters, given respectively by two sums on the right-hand side of Eq. 4, from the same four $\alpha_{s}$ scans with the fixed $\{1,1,1\}$ PDF as in Fig. 2. The curves from the four scans are stacked on the top of one another so as to fit in the figure. Stars indicate the minima of their corresponding parabolas. Positions of the stars for residuals can go as low as to $\alpha_{s}(M_{Z})=0.113$ for the $\{2,2,2\}$. The stars for the nuisance parameters are more stable between 0.118 and 0.120.
Figure 3: Separate contributions to $\chi^{2}_{R}/N_{\textrm{pt}}$ from the best-fit summed data residuals and nuisance parameters, as exemplified in Eq. 4, for the four ksys combinations shown in Fig. 2. The theory predictions are obtained with the CT25NNLO PDFs determined using ksys={1,1,1}.
@@
Figures 4 and 5 depict this dynamics in the complete $\alpha_{s}$ scans, now done using the CT25prel baseline data set that prefers a slightly lower $\alpha_{s}(M_{Z})<0.118$. In these figures, the PDFs are fitted for the indicated ksys and $\alpha_{s}(M_{Z})$.

### v2

@@ II.2 The CTEQ-TEA goodness-of-fit function
where $\chi^{2}_{R}$ quantifies the agreement of theory with experiments, and $\chi^{2}_{\textrm{LM}}$ is a generally small contribution imposing theoretical priors on the PDFs. The latter is constructed from Lagrange-Multiplier (LM) contributions whose role is to (i) enforce physical behavior of PDF flavor ratios in the extrapolation regions with poor constraints from the data and (ii) enable a “ridge” (L2) regularization on a few poorly constrained, yet correlated free parameters.
@@
$\chi^{2}_{E}(\bm{a},\bm{\lambda},\alpha_{s},\{m_{q}\})=$
$\sum_{k=1}^{N_{\textrm{pt}}^{(E)}}\frac{1}{s_{k}^{2}}\left(D_{k}-T_{k}\left(\bm{a},\alpha_{s},\{m_{q}\}\right)-\sum_{\alpha=1}^{N_{\lambda}^{(E)}}\beta_{k\alpha}\left(\bm{a},\alpha_{s},{m_{q}}\right)\lambda_{\alpha}\right)^{2}+\sum_{\alpha=1}^{N_{\lambda}^{(E)}}\lambda_{\alpha}^{2}\,.$ (3)
Here $D_{k}$ and $T_{k}$ are the $k$-th central experimental and theory values, respectively, $s_{k}=\sqrt{s^{2}_{k,\textrm{stat}}+s^{2}_{k,\textrm{uncor sys}}}$ is the total uncorrelated error on the measurement $D_{k}$, and $\beta_{k\alpha}^{(E)}$ is the correlation matrix of systematic uncertainties. $\bm{a}\equiv\{a_{1},a_{2}\dots,a_{N_{\textrm{pdf}}}\}$ and $\bm{\lambda}\equiv\{\lambda_{1},...\lambda_{N_{\lambda}^{(E)}}\}$ are the vectors of PDF parameters and systematic nuisance parameters of $E$, with the latter commonly assumed to be random and distributed according to the standard normal distribution $\mathcal{N}(0,1)$. Equation (3) indicates that $\chi^{2}$ generally depends on QCD parameters – $\alpha_{s}$ and quark masses $m_{q}$ – both through the theoretical predictions $T$ and the experimental correlation matrices $\beta_{k\alpha}$, although this dependence of the latter is usually neglected.
In our study, we minimize the total $\chi^{2}$ with respect to $\alpha_{s}$ and $\bm{a}$, while using constant $\{m_{q}\}$ and profiling [63] the nuisance parameters $\bm{\lambda}$ for every $\{\alpha_{s},\bm{a}\}$ combination. We will find it helpful that, at its global minimum for each $\alpha_{s}$, the $\chi^{2}$ can be expressed in terms of the best-fit PDF parameter combination $\bm{a}_{0}$ as (see Eq. (B8) in [16])
$\min_{\bm{a}}\chi^{2}(\bm{a},\alpha_{s},\{m_{q}\})\equiv\chi^{2}(\bm{a}_{0},\alpha_{s},\{m_{q}\})=\sum_{E}\left[\underbrace{\sum_{k=1}^{N_{\textrm{pt}}^{(E)}}[r_{0,k}^{(E)}]^{2}}_{\equiv D^{2}}+\underbrace{\sum_{\alpha=1}^{N_{\lambda}^{(E)}}[\lambda^{(E)}_{0,\alpha}]^{2}}_{\equiv R^{2}}\right],$ (4)
@@
and all terms on the right-hand side corresponding to experiment $E$. Furthermore, since $r_{0,k}^{(E)}$ and $\lambda^{(E)}_{0,\alpha}$ are correlated only through a relatively small number of PDF parameters, especially across separate experiments, in the first approximation they all can be assumed to be independently and normally distributed, $r_{0,k}^{(E)}\sim\mathcal{N}(0,1)$ and $\lambda^{(E)}_{0,\alpha}\sim\mathcal{N}(0,1)$. This independence opens the possibility of reclustering the individual data points in the total $\chi^{2}$ in Eq. (4), which will be pursued in Sec. IV. Furthermore, in Sec. IV.6 we can examine separate reclustering of sums of all best-fit residuals and nuisance parameters, denoted as $D^{2}$ and $R^{2}$, respectively.
This reclustering procedure does not affect the best fit but can be useful for understanding stochastic effects on the resultant uncertainty, especially in connection with a dynamic tolerance prescription.
II.3 Limitations of existing models for correlated systematics
Several aspects of the implementation of correlated systematic errors through $\beta_{k\alpha}$ in our $\chi^{2}$ introduced above will be relevant for interpreting the results that will follow.
@@ II.3 Limitations of existing models for correlated systematics
    We first note that the total number $N_{\lambda}$ of experimental nuisance parameters has grown from at most a few tens in the early global fits to more than 2000 in the CT25 analysis. That is, $N_{\lambda}$ is now close to a half of the total data points $N_{\textrm{pt}}$ due to the tendency of the latest experiments to provide large correlation or covariance matrices. In the best fit, many of these nuisance parameters are not normally distributed, but instead cluster close to zero, leading to a spike in the distribution.
@@
    Each nuisance parameter is included with a prior constraint, essentially an extra datapoint preventing the parameter’s completely arbitrary variations. Considerable attention has been paid recently to the fact that, by using the linear Gaussian models to publish their correlated systematic uncertainties, experiments unavoidably leave out critical information or introduce spurious assumptions that may proliferate when the number of priors is large. There are a couple such points, first, related to the standard assertion that all systematic nuisance parameters follow the standard normal distribution, $\lambda^{(E)}_{\alpha}\sim\mathcal{N}(0,1)$ in the $\chi^{2}$ definition in Eq. (3). Not only is it highly unlikely that all $\lambda_{\alpha}$ parameters will be thus distributed, but there are also known important cases where this assumption is false, e.g., for so-called two-point uncertainties that take the difference of at most two estimates, for instance, provided by two parton shower programs, as a measure of a possible uncertainty [64].
@@
    The second point of the information loss emerges from the common practice of publishing only normalized correlation matrices, $\sigma_{k\alpha}$, from which the actual correlation matrices, $\beta_{k\alpha}=X_{k}\sigma_{k\alpha}$ in Eq. (3), must be reconstructed by multiplying by unspecified reference values $X_{k}$ that are chosen within the PDF fit itself. The $X_{k}$ values are of order of either the central data values, $D_{k}$, or respective theoretical values, $T_{k}$. They do not need to coincide with either; the prescription for choosing $X_{k}$ is a delicate part of the fitting model [65, 66], approached differently by each PDF fitting group [67, 68], who try to reasonably fill the lacunas in the correlation models they get.
How do these considerations impact the precise determination of the QCD coupling constant? The next section offers some answers, after discussing another potential source of uncertainty due to the freedom in the PDF parametrization.
@@ III.1 The PDF parametrization uncertainty is small…
It has been pointed out over years, e.g., in  [15, 16, 69, 70], that QCD parameters determined from a global fit are mutually dependent on the PDFs and thus must be fitted simultaneously with the PDFs. Within the global fit, the correlation between the $\alpha_{s}$ and PDFs can be captured either by fitting them together or by scanning the $\chi^{2}(\alpha_{s},\mathbf{a}_{0}(\alpha_{s}))$ over $\alpha_{s}$, while minimizing the $\chi^{2}$ with respect to the PDFs for each $\alpha_{s}$ value. Either approach can be used to determine the best-fit combination of $\alpha_{s}$ and PDFs together with their uncertainties. We will employ both, especially the latter one – the $\alpha_{s}$ scan.
Perhaps unexpectedly, our study found a low correlation between the best-fit $\alpha_{s}$ and the functional forms of PDF parametrizations in the vicinity of the overall global $\chi^{2}$ minimum. The consequence is that the choice of the PDF parametrizations has a low impact on the $\alpha_{s}$ and its error, among the candidate functional forms of the CT25 fit that render the lowest achieved values of $\chi^{2}$. The variation of the best-fit $\alpha_{s}$ caused by the choice of the parametrization forms is comparable to the uncertainty corresponding to a $\Delta\chi^{2}=1$ change in the global $\chi^{2}$. We will see that this is quite a small uncertainty compared to other sources.
@@
Figure 1 illustrates the low correlation of best-fit $\alpha_{s}$ and PDFs by comparing the $\chi^{2}$ dependence from an $\alpha_{s}$ scan for the functional forms of the nominal CT25 fit and 287 fits with alternative PDF parametrizations explored as a part of the full CT25 exercise. In this plot, we focus on the component of the total $\chi^{2}$ consisting from experimental log-likelihoods and denoted by $\chi^{2}_{R}$ as in Eq. (1), and neglecting the prior component $\chi^{2}_{\textrm{LM}}$ that is relatively constant in this comparison. All fits in this figure are done with identical settings that are close both to the CT25prel and final CT25 ones introduced in Sec. II.1. Namely, the shown fit with the nominal CT25 parametrization achieves $\min\chi^{2}_{R}\equiv\chi^{2}_{0}=5087$ for $N_{\textrm{pt}}=4260$ data points. Compared to the CT25prel fit, it excludes the CDHSW heavy-nucleus CC DIS data set [34], which prefers a low $\alpha_{s}$, as well as the weakly sensitive H1 $F_{L}$ data set [38]. It retains other experiments of the CT25prel baseline.
The orange star in Fig. 1 indicates the best-fit combination $\{\alpha_{s}\approx 0.1185,\chi^{2}_{R}=5087\}$ obtained with the nominal CT25 parametrization. The orange curve indicates the $\chi^{2}_{R}$ dependence from the $\alpha_{s}$ scan with this parametrization. This curve closely follows the parabolic behavior. The vertical magenta band indicates the interval of $\alpha_{s}$ values corresponding to the increase $\Delta\chi^{2}_{R}\leq 1$ from the $\alpha_{s}$ scan, measured with respect to the minimal $\chi^{2}_{R}$ on this quasi-parabola.
@@
Here we observe a striking behavior: while the $\chi^{2}$ variations among the candidate fits are sizable in the vertical ($\chi^{2}$) direction, they have modest horizontal variations, compatible with the size of the $\Delta\chi^{2}=1$ interval, as far as the best-fit $\alpha_{s}$ are concerned. The picture immediately reminds of the textbook case of the $\chi^{2}$ variability due to approximately Gaussian random fluctuations in data [71]: when one plots the log-likelihood ($\chi^{2}$) vs. the parameter of interest, e.g., $\alpha_{s}$, vertical fluctuations of the resulting parabola fall into the interval $\chi^{2}_{R}\leq\chi^{2}_{0}+\sqrt{2N_{\textrm{pt}}}$ in $\approx 68\%$ of cases, while about 68% of the horizontal fluctuations fall within the interval $\Delta\chi^{2}_{R}\leq 1$.
In our example, the best-fit $\alpha_{s}$ from fits with alternative parametrizations fall within the $\Delta\chi^{2}_{R}=1$ interval from the scan, reflecting statistical fluctuations in the data at 68% CL for the fixed data set and PDF functional forms. This interval corresponds to the uncertainty $\delta\alpha_{s}\approx\pm 0.0005$, and this conclusion holds for other candidate selections of baseline experiments. Namely, while the preferred $\alpha_{s}$ value moves from $\approx 0.1177$ for the CT25prel to 0.1185 for the CT25 selection of data sets, see the bottom of Table 1, the pattern of the parametrization dependence is about the same in these fits as in Fig. 1, with the parametrization uncertainty staying at the level of 0.0005.
The conclusion based on the ensemble of nearly 300 PDF parametrizations is then that they give very similar $\alpha_{s}$ values, showing that the determination of $\alpha_{s}$ is essentially independent of the chosen PDF parametrization. If only this source of uncertainty were considered, the uncertainty span of approximately $\pm 0.0005$ about the central value would be smaller than both the PDG uncertainty [72] of $\pm 0.0009$ and errors of order $\pm 0.002$ quoted by dedicated PDF-only fits [73, 17, 74, 20, 19, 16, 75]. Given these observations, the final $\alpha_{s}$ uncertainty estimate does not include a component arising from PDF parametrizations.
@@
The global fit contains two additional notable sources of $\alpha_{s}$ uncertainties: (1) the dependence on the models of correlated systematic errors discussed in Sec. II.3 and (2) the tensions between preferred values of $\alpha_{s}$ from the separate categories of data sets. This section addresses the first aspect and illustrates the possible impact of the choice of the reference normalization factors $X_{k}$ in the experimental correlation matrices $\beta_{k\alpha}=X_{k}\sigma_{k\alpha}$, which is controlled in our fit by the parameter ksys. Section IV will follow with addressing the second aspect.
Since the experiments normally do not publish their recommended $X_{k}$, a variety of “guesses” about $X_{k}$ can be, and have been, made in the past PDF fits [65, 66]. As illustrative possibilities, we consider two such prescriptions, which we label as “ksys=1” and “ksys=2”. According to the ksys=1 prescription, each correlated systematic element $\beta_{k\alpha}$ (initially provided as a percentage fraction $\sigma_{k\alpha}$) is normalized to the current theory prediction $T_{k}$ for each data point. This treatment is more suitable when the corresponding experimental error is multiplicative. Major systematic errors for the LHC experiments, for example jet energy scale systematic errors, are treated as multiplicative by the experiments. [^4] The luminosity error of LHC inclusive jet cross sections is treated as additive. On the other hand, this prescription can introduce a theory bias in the effect of systematic errors, reflected e.g. in the dependence on the version of the “tt” method for normalization of correlated errors seen in the past analyses [68]. The majority of fits by CTEQ-TEA and other groups choose ksys=1 for all correlated errors as a default.
@@
For the ensuing discussion, we divide our baseline data set of measurements into three subsets: 1) deep-inelastic scattering (DIS), 2) Drell-Yan (DY) pair production, 3) inclusive jet and $t\bar{t}$ production. In each subset, we normalize all $\beta_{k\alpha}$ as though they are multiplicative (ksys=1) or additive (ksys=2), with the caveat that the ksys=1 option is more likely to be closer to the truth. Measurements do contain a mix of multiplicative and additive errors. The true value of $\alpha_{s}$ extracted from the global fit is somewhere between the extremes corresponding to the “all ksys=1” and “all ksys=2” options, but likely closer to ksys=1. This setup allows us to explore sensitivity to systematic error modeling, particularly for hadron jet production observables that exhibit a very large sensitivity.
Impact of systematic-error normalizations for fixed PDFs.
We first demonstrate what happens in the final CT25 fit when we change the default setting of using “ksys=1” in all three measurement categories, labeled as $\{\texttt{ksys}_{DIS},\texttt{ksys}_{DY},\texttt{ksys}_{\textrm{Jet}+t{\bar{t}}}\}$ = $\{1,1,1\}$. First, to illustrate how the ksys choice modifies $\chi^{2}$ when all other settings are unchanged, let us examine the $\chi^{2}$ dependence for several combinations of ksys values while using the same PDF ensemble. Figure 2 illustrates these changes, by showing $\chi^{2}$ vs $\alpha_{s}$ computed with $\{1,1,1\}$, $\{1,1,2\}$, $\{1,2,2\}$, and $\{2,2,2\}$ ksys choices, each using the same nominal CT25 data set and best-fit PDFs obtained with $\{1,1,1\}$.
@@
It is interesting to investigate further into the details of the fits to understand the driving mechanism of these shifts. First, it can be noted that, as far as the $\alpha_{s}$ dependence goes, the data-point residuals are more responsive to the form of the systematic errors than the best-fit nuisance parameters. This can be seen from Fig. 3 that separately plots the contributions from the best-fit residuals and nuisance parameters, given respectively by two sums on the right-hand side of Eq. (4), from the same four $\alpha_{s}$ scans with the fixed $\{1,1,1\}$ PDF as in Fig. 2. The curves from the four scans are stacked on top of one another so as to fit in the figure. Stars indicate the minima of their corresponding parabolas. Positions of the stars for residuals can go as low as $\alpha_{s}(M_{Z})=0.109$ for $\{2,2,2\}$. The stars for the nuisance parameters stay between 0.117 and 0.120.
Figure 3: Separate contributions to $\chi^{2}_{R}/N_{\textrm{pt}}$ from the best-fit summed data residuals and nuisance parameters, as exemplified in Eq. (4), for the four ksys combinations shown in Fig. 2. The theory predictions are obtained with the CT25NNLO PDFs determined using ksys={1,1,1}.
@@
Impact of systematic-error normalizations with fitted PDFs.
Figures 4 and 5 depict this dynamics in the complete $\alpha_{s}$ scans, now done using the CT25prel baseline data set that prefers a slightly lower $\alpha_{s}(M_{Z})<0.118$. In these figures, the PDFs are fitted for the indicated ksys and $\alpha_{s}$.